OpenStudy (breet1718):

PLEASE HELP ME!! Which of the following is a solution to the equation c + (4 – 3c) – 2 = 0?

TheSmartOne (thesmartone):

c + (4 – 3c) – 2 = 0 Let's get rid of the parenthesis. Since we're adding, we can just get rid of it. If we were subtracting the value in the parenthesis, we would need to distribute the negative inside the parenthesis to remove it. But since we're adding, distributing the positive changes nothing. c + 4 – 3c – 2 = 0 Do you know about like terms or solving for a variable?

OpenStudy (breet1718):

I think so but im not sure im a little rusty on it

TheSmartOne (thesmartone):

c + 4 – 3c – 2 = 0 What do you get when you do c - 3c? c - 3c = ??

OpenStudy (breet1718):

4c?

TheSmartOne (thesmartone):

Not c + 3c c - 3c If we factored the c out, we would have c(1 - 3) what is 1 - 3 = ?

OpenStudy (breet1718):

-2?

TheSmartOne (thesmartone):

yup, and c(1-3) = c(-2) = -2c ab is the same thing as ba but we like to keep the variable after the number so now we have -2c + 4 - 2 = 0 what is 4 - 2 = ?

OpenStudy (breet1718):

2

TheSmartOne (thesmartone):

now we have -2c + 2 = 0 add 2c to both sides, what do you get?

OpenStudy (breet1718):

2=0+2c?

TheSmartOne (thesmartone):

yes and 0 + 2c is nothing but 2c we now have 2 = 2c what do you get if you divide 2 on both sides?

OpenStudy (breet1718):

1=c?

TheSmartOne (thesmartone):

and that's your answer :)

OpenStudy (breet1718):

OMG! thank you so much! i was stuck on that question for an hour! now it makes sense!

TheSmartOne (thesmartone):

Anytime! Welcome to OpenStudy! :D

OpenStudy (breet1718):

thank you!